OpenStudy (anonymous):

13. Find the center and radius of the circle with equation (x-5)^2+(y+6)^2=9

OpenStudy (turingtest):

a circle with center \((h,k)\) has the equation\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

Where r-->radius of circle

OpenStudy (anonymous):

Can you follow the instruction?

OpenStudy (anonymous):

as turingtest wrote , match up \[(x-5)^2+(y+6)^2=9\] with \[(x-h)^2+(y-k)^2=r^2\] identify \(h\) and \(k\) and \(r\)

OpenStudy (anonymous):

r = 3 and the center is (5,-6)

OpenStudy (turingtest):

yes :)